A bus is moving with a= 5m/s2 from rest find its velocity after2sec.


1
Expert's answer
2020-10-12T07:50:44-0400

solution

given data

initial velocity(u)=0 m/s

acceleration (a)=5 m/s^2


By using the first equation of motion

"v=u+at"

at t=2 sec it's velocity will be

"v=0+5\times2=10 m/s"


velocity after 2 second will be 10 m/s.